OpenStudy (smartsass):

How do u use a ti-84 plus to solve cramer's rule

OpenStudy (ddcamp):

Do you specifically need to use Cramer's rule, or are you just solving systems on your TI?

OpenStudy (smartsass):

Solving systems on ti.

OpenStudy (anonymous):

i have one 2, but i cant do anything but the basic stuff on it... :(

OpenStudy (smartsass):

Ok thxs!

OpenStudy (ddcamp):

To just solve a system, the easiest thing to use is reduced row echelon form (rref) To use this, create a matrix. If you have 2 equations, you want a 2×3 matrix, if you have 3 equations, you want a 3×4 matrix, and so on: [2nd] [Matrix] [→] [→] [Enter] Select the dimensions you want, then put in the coefficients. If your equations are: \[3a + 2b = 9 \\ a - 2b = 11\] Your matrix will look like: \[\left[\begin{matrix}3 & 2 & 9 \\ 1 & -2 & 11 \end{matrix}\right]\] Now, you have your matrix (and it should automatically be saved as [A]). To solve the equation, input rref([A]). Find rref( under the matrix math menu: [2nd] [Matrix] [→] Then matrix [A] : [2nd] [Matrix] [A] The output of the above matrix is: \[\left[\begin{matrix}1 & 0 & 5 \\ 0 & 1 & -3\end{matrix}\right]\] Which means that a=5 and b=-3. A bit of work to learn how to do it on the calculator, but this can be used for much larger systems as well.

OpenStudy (smartsass):

Thank you so much, I really appreciate it!